The supplement of an angle measures 40" more than twice its complement. Find the measure of the angle.
Answer by ankor@dixie-net.com(22740)

Assume you mean 40 degrees, not 40 seconds
The supplement of an angle measures 40 degrees more than twice its complement.
Find the measure of the angle.
:
Let x = the angle
then
(90-x) = it's complement
and
(180-x) = it's supplement
:
Write an equation for the statement:
"The supplement of an angle measures 40 degrees more than twice its complement."
180 - x = 2(90-x) + 40
:
180 - x = 180 - 2x + 40
:
-x + 2x = 180 - 180 + 40
:
x = 40 degrees
:
:
Check solution in the statement
The supplement of an angle measures 40 degrees more than twice its complement.
180 - 40 = 2(90-40) + 40
140 = 2(50) + 40